OpenStudy (anonymous):

I need this answered really soon, so if you guys can do that for that would be great! :) Thanx! LUV U! Simplify. 8 – (–4)(–3)

OpenStudy (aravindg):

Ok :) So lets start with second term -4*-3=?

OpenStudy (anonymous):

er whut?

OpenStudy (aravindg):

Remember product of 2 negative numbers always gives a positive number.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2 like terms = positive

OpenStudy (aravindg):

\[-4 \times -3=\]

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

yeah, thats not in my answers...

OpenStudy (aravindg):

We are doing multiplication of -4 and -3.

OpenStudy (aravindg):

what did u get when you multiplied -4 and -3.

OpenStudy (aravindg):

?

OpenStudy (anonymous):

-7.

OpenStudy (aravindg):

No!!

OpenStudy (aravindg):

4 times 3 means 3+3+3+3 and not 4+3

OpenStudy (anonymous):

ug this isnt making any sense

OpenStudy (aravindg):

Why not?

OpenStudy (anonymous):

idk... I aint no teacher, the only reason i came on the web site is cuz i needed help in math...

OpenStudy (anonymous):

And im not getting it

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

i gotta go... thax 4 trying tho... :) I appreciate it alot more than you think. :) :/

OpenStudy (aravindg):

Oh sorry you didnt get the help. If you stay I can try again?

OpenStudy (anonymous):

okay. :) thanx Urg r u a teacher?

OpenStudy (aravindg):

Nope I am just a student :)

OpenStudy (anonymous):

what grade?

OpenStudy (aravindg):

I am in college.

OpenStudy (anonymous):

Oh good! I was hoping sum 1 older than me! :)

OpenStudy (anonymous):

im in 7th

OpenStudy (anonymous):

okay lets start... shmurr

OpenStudy (anonymous):

im in k12 so yeah... all homework

OpenStudy (aravindg):

Good. So lets get back to your question (4)(1)=4 (4)(2)=4+4=8 (4)(3)=4+4+4=12 ANY PROB?

OpenStudy (anonymous):

im sorry, but how did we get here??

OpenStudy (anonymous):

how did we get to the thing u just posted?

OpenStudy (anonymous):

Iv gotta know every step

OpenStudy (aravindg):

Any number times any number means the number of times it is added 2*2 means 2 added 2 times 5*3 means 5+5+5 4*3 means 4+4+4

OpenStudy (anonymous):

good so 12

OpenStudy (aravindg):

great

OpenStudy (anonymous):

okay, thanx! Il tell you how i do on my assessment! <3 (in like 5 min)

OpenStudy (aravindg):

Good :) ..Hope you do well! All the best!

OpenStudy (anonymous):

alrghty! The answer is!!!: 100% boo ya. :) lol

OpenStudy (aravindg):

Yey!! :D Good work!!!

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

Okay, My bro and sister have to work too so I dont want to take up too much bandwidth, so Im gonna leave ths site and they can work on their laptops cuz theres only 2 bars on the wifi thing so c ya! and thanx! <3

OpenStudy (anonymous):

il be back! lol um yeah cya

OpenStudy (aravindg):

Bye! And do visit again..There are many others like me who can help you. Seeya!